Grade 12th passGeneral Physics

10ml of co is mixed with 25ml of air having 20% 02 by the volume .what would be the final volume if none of co and 02 after left the reaction

Deepjyoti mech
9 Years agoGrade 12th pass

1 Answer

Vikas TU
9 Years ago
Dear Student,
 
25 ml of air has 20% O2 by volume
Amount of oxygen present = 20/100 *25 = 5 ml
If none  of CO and O2 is left after the reaction then the final volume will be 10+5 = 15ml
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)